# Stochastic Calculus for Pathwise Observables of Markov-Jump Processes: Unification of Diffusion and Jump Dynamics

### v1

@@ IV.3 Thermodynamic Transport Bound
where $\bm{H}(\tau)]_{ij}=z_{i}(\tau)-z_{j}(\tau)$ is the difference of some observable state function $z_{\tau}\equiv z(x_{\tau},\tau)=\sum_{k}\delta_{k,x_{\tau}}z_{k}(\tau)$. The mean of $B_{t}$ vanishes while the second moment reads
@@
which we can identify with the short-time fluctuation function $\mathcal{D}_{t}=\langle B^{2}_{t}\rangle/t$. Namely, consider the infinitesimal difference $\mathrm{d}z(\tau)=z_{\tau+\mathrm{d}\tau}-z_{\tau}$ along the trajectory. Then for any $n\in\mathds{N}_{>0}$ we then get $\langle[\mathrm{d}z(\tau)]^{n}\rangle=\sum_{x,y}(z_{y}-z_{x})^{n}r_{xy}p_{x}(\tau)\mathrm{d}\tau+\mathcal{O}(\mathrm{d}\tau^{2})$. Thus, $\mathrm{var}[\mathrm{d}z(\tau)]=\sum_{x,y}(z_{y}-z_{x})^{2}r_{xy}p_{x}(\tau)\mathrm{d}\tau+\mathcal{O}(\mathrm{d}\tau^{2})$ and
@@
The average of $z_{t}$ (and of any vector) is given by $\langle z_{\tau}\rangle=\sum_{k}\sum_{l}\delta_{k,l}z_{k}(\tau)p_{l}(t)$. By combining the various averages entering the Cauchy-Schwarz inequality $\langle A_{t}B_{t}\rangle\leq\langle A_{t}^{2}\rangle\langle B_{t}^{2}\rangle$ we obtain the transport bound for MJP
@@ IV.3.1 Saturation of the Transport Bound
be the generator for a fully connected three state MJP with states $\{1,2,3\}$. Saturation of Eq. (93) entails choosing $z_{\tau}$ such that $\bm{H}(\tau)^{T}\propto z_{\tau}$ to saturate the Cauchy-Schwarz inequality. Explicitly, for all $\tau$
@@ IV.3.2 Unification of (C)TUR and TB
Equation (99) therefore naturally relates the TB to the CTUR as a special case. Indeed, the gTB is efficient, i.e. yields a non-zero bound in both, transient and stationary systems, while the TB only yields nontrivial bounds in transient systems. A similar connection can be made in continuous space, where a recent TUR for underdamped Langevin dynamics [126] can be seen as a generalization of the continuous TB [51]. It should be noted that, while the TB in general cannot be saturated, the gTB can be saturated with the choice of weight $\bm{\kappa}(\tau)\propto z_{\tau}$. Moreover, in some cases the transient TB of one system may be identical to the stationary gTB of another system. The following example, an extension of the SAT model (see Fig. 6b), shows how the TB may give rise to the gTB in a special case.
@@
The SAT describes the transport of a large molecule (denoted molecule $A$) through a cell membrane against its chemical gradient using the transport of helper molecule (denoted molecule $B$) [145, 146, 147], see Fig. 6a-b. The reduction to a six-state Markov model, $\mathcal{N}=\{1,2,3,4,5,6\}$, is commonly used to describe this secondary active transport [145, 146, 147]. The efficiency of the process is related to the number of molecules $A$ transported given a concentration of $B$. The transport can be split into two categories: symporters and antiporters. These refer to molecule $A$ being transported parallel or anti-parallel to molecule $B$, respectively. To effectively describe the total number of transported molecules, we extend the six-state model to a chain of six-state models, i.e., $\tilde{\mathcal{N}}=\mathcal{N}\times\mathds{Z}$, where the additional degree of freedom corresponds to the number of $A$ molecules transported through the membrane. We are interested in the fraction of the total steady-state entropy production that we can estimate via the thermodynamic bounds for various values of transition rates ${\rm e}^{\mathrm{out}}_{B}$. These transition rates are assumed proportional to the concentration of $B$ outside the cell [145].
To make reasonable comparisons, we either have to consider symporter or antiporter systems. We are going to focus on the latter. We first vary ${\rm e}^{\mathrm{out}}_{B}$ to assess which values of transition rates give rise to antiporter behaviour before adressing thermodynamic inferrence.
@@
However, the current (entering the stationary TUR) in the six-state model gives the insight we require. The relevant transition weight is $\kappa_{ij}^{\mathrm{SAT}}=\delta_{i1}\delta_{j3}-\delta_{i3}\delta_{j1}$, effectively counting the total number of molecules $A$ being transported towards the exterior of the cell. The mean current is $\langle J_{t}^{\mathrm{SAT}}\rangle_{\mathrm{s}}=\sum_{x,y}\kappa_{xy}^{\mathrm{SAT}}r_{xy}p_{x}^{\mathrm{s}}$. If $\langle J_{t}^{\mathrm{SAT}}\rangle_{\mathrm{s}}>0$, the average total number of molecules $A$ transported towards the outside of the cell is positive, corresponding to an average transport of $A$ from the inside to the outside of the cell. Thus, the condition for an antiporter behavior is $l_{A}/e^{\mathrm{out}}_{A}>p_{3}^{\mathrm{s}}/p_{1}^{\mathrm{s}}$, which depends on ${\rm e}^{\mathrm{out}}_{B}$ via the steady-state distribution. With the parameters we use, see Tab. 3, we need $p_{3}^{\mathrm{s}}/p_{1}^{\mathrm{s}}<2/3$. The value $2/3$ is marked as a black line in the inset of Fig. 9a, where the green line corresponds to the value of $p_{3}^{\mathrm{s}}/p_{1}^{\mathrm{s}}$ as a function of ${\rm e}^{\mathrm{out}}_{B}$. The dashed red line marks the value of ${\rm e}^{\mathrm{out}}_{B}$ s.t. $\mathcal{A}_{\mathcal{C}_{3}}=0$, while the dashed orange line is where $p_{3}^{\mathrm{s}}/p_{1}^{\mathrm{s}}=2/3$. The same can be readily seen in Fig. 9b, where the mean current is presented for various values of ${\rm e}^{\mathrm{out}}_{B}$. Clearly, the condition $\mathcal{A}_{\mathcal{C}_{3}}=0$ is not sufficient for a positive current.
Additionally, we can consider the transient extended model and apply the TB. For the extended model, consider $p_{x,n}(\tau)$ the probability of being in state $x$ in system $n$ at time $\tau$, where $n\in\mathds{Z}$ corresponds to the number of $A$ molecules transported through the membrane. When marginalizing over $\mathds{Z}$ we recover the steady-state probability $p_{x}^{\mathrm{s}}=\sum_{n\in\mathds{Z}}p_{x,n}(\tau)$. As a state function we choose the number of $A$ molecules $z_{\tau}=n$, such that
$\langle z_{\tau}\rangle=\sum_{n\in\mathds{Z}}n\sum_{x=1}^{6}p_{x,n}(\tau)=\sum_{n\in\mathds{Z}}n\tilde{p}_{n}(\tau),$ (103)
@@
$\mathcal{D}_{t}^{\mathrm{ext}}=\frac{1}{t}\sum_{n\in\mathds{Z}}\sum_{x,y\in\mathcal{N}}\int_{0}^{t}\mathrm{d}\tau$ $\left[(z_{y,n}-z_{x,n})^{2}\delta_{x\to y,n}+(z_{y,n+1}-z_{x,n})^{2}\delta_{x\to y,n\to n+1}\right$.
@@
$\mathcal{D}_{t}^{\mathrm{ext}}=\frac{1}{t}\sum_{n\in\mathds{Z}}\int_{0}^{t}\mathrm{d}\tau\left[r_{13}p_{1,n}(\tau)+r_{31}p_{3,n}(\tau)\right],$ (105)
@@
The quality factors of the TUR $\mathcal{Q}^{\mathrm{TUR}}$ and TB $\mathcal{Q}^{\mathrm{TB}}$ are shown in Fig. 9c (solid and dashed line, respectively) for ${\rm e}^{\mathrm{out}}_{B}$ in the antiporter regime. Because $t\mathcal{D}_{t}^{\mathrm{ext}}>\mathrm{var}_{\mathrm{s}}(J_{t}^{\mathrm{SAT}})$ for all $t>0$, the TUR infers a higher fraction of the entropy production, i.e., $\mathcal{Q}^{\mathrm{TUR}}\geq\mathcal{Q}^{\mathrm{TB}}$. Additionally, one may show that taking ${\rm e}^{\mathrm{out}}_{B}\to\infty$ leads to the quality factors converging to a finite value [12, 152]. Hence, there is no improvement in the inference of dissipation at higher concentrations of $B$ outside the cell beyond a certain point.
@@ IV.4 Thermodynamic Correlation Bound
where $\tilde{Z}_{xy}(\tau)=Z_{xy}(\tau)\left[z_{x}(\tau)+z_{y}(\tau)-2F(\tau)\right]$ for some time-dependent function $F(\tau)$. The second key ingredient is the shifted and rescaled density observable
${C}_{t}$ $=\frac{1}{\sqrt{t}}\int_{\tau=0}^{\tau=t}\left[\bm{V}-\langle{V}_{\tau}\rangle\bm{1}\right]^{T}\mathrm{d}\bm{\tau}(\tau),$ (108)
where $\bm{1}$ is the one-vector [i.e. $(\bm{1})_{i}=1$ for all $i\in\mathcal{N}$] and $\langle V_{\tau}\rangle=\sum_{i}V_{i}p_{i}(\tau)$. While the TUR and TB allow the state functions $z_{\tau}$ and $\bm{V}(\tau)$ to be explicitly time-dependent, the correlation bound requires these to be strictly time-independent. Furthermore, we write, e.g., $\langle z_{\tau}\rangle=\sum_{i}z_{i}p_{i}(\tau)$ to explicitly highlight that the average depends on time through $p_{i}(\tau)$. Additionally, we need to introduce the re-weighted two-point probability
@@
With the shorthand notations $\bm{V}^{\Delta}(\tau)=\bm{V}-\langle{V}(\tau)\rangle\bm{1}$ and $z^{F}_{\tau}=z_{\tau}-\langle F(\tau)\rangle\bm{1}$, the cross term is computed to be
$t\langle B_{t}C_{t}\rangle$ $=\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{x,y,i}\mathds{1}_{\tau<\tau^{\prime}}V^{\Delta}_{i}(\tau^{\prime})\left(H_{yx}-Z_{xy}(\tau)\left[z_{x}^{F}(\tau)+z_{y}^{F}(\tau)\right]r_{xy}p_{x}(\tau)\right)\left[P(i,\tau^{\prime}|y,\tau)-P(i,\tau^{\prime}|x,\tau)\right],$ (113)
@@
$\sum_{x,y}Z_{xy}(z_{x}^{F}+z_{y}^{F})r_{xy}p_{x}\left[P(i|y)-P(i|x)\right]$
$=\sum_{x,y}z_{x}^{F}(r_{xy}p_{x}-r_{yx}p_{y})\left[P(i|y)-P(i|x)\right].$ (115)
Using time-homogeneity of the semi-group $\bm{P}(\tau^{\prime}|\tau)=\bm{P}(\tau^{\prime}-\tau)\equiv{\rm e}^{\bm{L}(\tau^{\prime}-\tau)}$ which obeys the master equation Eq. (32), i.e. $\partial_{\tau}\bm{P}(\tau)=\bm{L}\bm{P}(\tau)=\bm{P}(\tau)\bm{L}$, and $\partial_{\tau}\bm{P}(\tau^{\prime}-\tau)=-\partial_{\tau^{\prime}}\bm{P}(\tau^{\prime}-\tau)$, the integrand in Eq. (113) simplifies to $2\sum_{x,i}\mathds{1}_{\tau<\tau^{\prime}}V^{\Delta}_{i}(\tau^{\prime})z^{F}_{x}(\tau)p_{x}(\tau)\partial_{\tau^{\prime}}P(i,\tau^{\prime}|x,\tau)$. Since
@@
$t\langle B_{t}C_{t}\rangle$ $=2\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\mathds{1}_{\tau<\tau^{\prime}}\partial_{\tau^{\prime}}\mathrm{cov}({V}_{\tau^{\prime}},z_{\tau}).$ (117)
An integration by parts in $\tau^{\prime}$ with $\partial_{\tau^{\prime}}\mathds{1}_{\tau<\tau^{\prime}}=\delta(\tau^{\prime}-\tau)$ further yields
@@
$\leq\frac{\langle A_{t}^{2}\rangle_{\mathrm{s}}}{t}\mathrm{pvar}_{\mathrm{ps}}^{F}({z})\leq\frac{\Delta S_{\mathrm{tot}}(t)}{2t}\mathrm{pvar}_{\mathrm{ps}}^{F}({z}).$ (123)
@@
Example II: 4-State Toy model.—Consider a four-state model $\{1,2,3,4\}$ on a ring, see Fig. 6e. The rates clockwise are unity and the counterclockwise rates are controlled by the parameter $\epsilon$. The steady-state distribution $p_{i}^{\mathrm{s}}=p=1/4$ is independent of $\epsilon$. Suppose that we can only distinguish between a pair of mesostates A and B it can only be distinguished whether the states are in either states 1 and 2 or states 3 and 4. These we refer to mesostates A (containing $\{1,2\}$) and B (containing $\{3,4\}$), see Fig. 6e.

### v2

@@ IV.3 Thermodynamic Transport Bound
where $\bm{H}(\tau)]_{ij}=z_{i}(\tau)-z_{j}(\tau)$ is the difference of some observable state function $z_{\tau}\equiv z(x_{\tau},\tau)=\sum_{k}{\color[rgb]{0,0,0}\delta_{x_{\tau}k}}z_{k}(\tau)$. The mean of $B_{t}$ vanishes while the second moment reads
@@
which we can identify with the short-time fluctuation function $\mathcal{D}_{t}=\langle B^{2}_{t}\rangle/t$. Namely, consider the infinitesimal difference $\mathrm{d}z(\tau)=z_{\tau+\mathrm{d}\tau}-z_{\tau}$ along the trajectory. Then for any $n\in{\color[rgb]{0,0,0}\mathbb{N}}_{>0}$ we then get $\langle[\mathrm{d}z(\tau)]^{n}\rangle=\sum_{x,y}(z_{y}-z_{x})^{n}r_{xy}p_{x}(\tau)\mathrm{d}\tau+\mathcal{O}(\mathrm{d}\tau^{2})$. Thus, $\mathrm{var}[\mathrm{d}z(\tau)]=\sum_{x,y}(z_{y}-z_{x})^{2}r_{xy}p_{x}(\tau)\mathrm{d}\tau+\mathcal{O}(\mathrm{d}\tau^{2})$ and
@@
The average of $z_{t}$ is given by $\langle z_{\tau}\rangle=\sum_{k}\sum_{l}\delta_{kl}z_{k}(\tau)p_{l}(t)$. By combining the various averages entering the Cauchy-Schwarz inequality $\langle A_{t}B_{t}\rangle\leq\langle A_{t}^{2}\rangle\langle B_{t}^{2}\rangle$ we obtain the transport bound for MJP
@@ IV.3.1 Saturation of the Transport Bound
be the generator for a fully connected three state MJP with states $\{1,2,3\}$. Saturation of Eq. (101) entails choosing $z_{\tau}$ such that $\bm{H}(\tau)^{T}\propto{\color[rgb]{0,0,0}\bm{Z}(\tau)}$ to saturate the Cauchy-Schwarz inequality. Explicitly, for all $\tau$
@@ IV.3.2 Unification of (C)TUR and TB
Equation (107) therefore naturally relates the TB to the CTUR as a special case. Indeed, the gTB is efficient, i.e. yields a non-zero bound in both, transient and stationary systems, while the TB only yields nontrivial bounds in transient systems. A similar connection can be made in continuous space, where a recent TUR for underdamped Langevin dynamics [104] can be seen as a generalization of the continuous TB [40]. It should be noted that, while the TB in general cannot be saturated, the gTB can be saturated with the choice of weight $\bm{\kappa}(\tau)\propto{\color[rgb]{0,0,0}\bm{Z}(\tau)}$. Moreover, in some cases the transient TB of one system may be identical to the stationary gTB of another system. The following example, an extension of the SAT model (see Fig. 6b), shows how the TB may give rise to the gTB in a special case.
@@
To make reasonable comparisons in the SAT model, we either have to consider symporter or antiporter systems. We are going to focus on the latter. We first vary ${\color[rgb]{0,0,0}e}^{\mathrm{out}}_{B}$ to assess which values of transition rates give rise to antiporter behaviour before adressing thermodynamic inferrence.
@@
However, the current (entering the stationary TUR) in the six-state model gives the insight we require. The relevant transition weight is $\kappa_{ij}^{\mathrm{SAT}}=\delta_{i1}\delta_{j3}-\delta_{i3}\delta_{j1}$, effectively counting the total number of molecules $A$ being transported towards the exterior of the cell. The mean current is $\langle J_{t}^{\mathrm{SAT}}\rangle_{\mathrm{s}}=\sum_{x,y}\kappa_{xy}^{\mathrm{SAT}}r_{xy}p_{x}^{\mathrm{s}}$. If $\langle J_{t}^{\mathrm{SAT}}\rangle_{\mathrm{s}}>0$, the average total number of molecules $A$ transported towards the outside of the cell is positive, corresponding to an average transport of $A$ from the inside to the outside of the cell. Thus, the condition for an antiporter behavior is $l_{A}/e^{\mathrm{out}}_{A}>p_{3}^{\mathrm{s}}/p_{1}^{\mathrm{s}}$, which depends on ${\color[rgb]{0,0,0}e}^{\mathrm{out}}_{B}$ via the steady-state distribution. With the parameters we use, see Tab. 4, we need $p_{3}^{\mathrm{s}}/p_{1}^{\mathrm{s}}<2/3$. The value $2/3$ is marked as a black line in the inset of Fig. 9a, where the green line corresponds to the value of $p_{3}^{\mathrm{s}}/p_{1}^{\mathrm{s}}$ as a function of ${\color[rgb]{0,0,0}e}^{\mathrm{out}}_{B}$. The dashed red line marks the value of ${\color[rgb]{0,0,0}e}^{\mathrm{out}}_{B}$ s.t. $\mathcal{A}_{\mathcal{C}_{3}}=0$, while the dashed orange line is where $p_{3}^{\mathrm{s}}/p_{1}^{\mathrm{s}}=2/3$. The same can be readily seen in Fig. 9b, where the mean current is presented for various values of ${\color[rgb]{0,0,0}e}^{\mathrm{out}}_{B}$. Clearly, the condition $\mathcal{A}_{\mathcal{C}_{3}}=0$ is not sufficient for a positive current.
[…]
$\langle z_{\tau}\rangle=\sum_{n\in{\color[rgb]{0,0,0}\mathbb{Z}}}n\sum_{x=1}^{6}p_{x,n}(\tau)=\sum_{n\in{\color[rgb]{0,0,0}\mathbb{Z}}}n\tilde{p}_{n}(\tau),$ (111)
@@
$\mathcal{D}_{t}^{\mathrm{ext}}=\frac{1}{t}\sum_{n\in{\color[rgb]{0,0,0}\mathbb{Z}}}\sum_{x,y\in\mathcal{N}}\int_{0}^{t}\mathrm{d}\tau$ $\left[(z_{y,n}-z_{x,n})^{2}\delta_{x\to y,n}+(z_{y,n+1}-z_{x,n})^{2}\delta_{x\to y,n\to n+1}\right$.
@@
$\mathcal{D}_{t}^{\mathrm{ext}}=\frac{1}{t}\sum_{n\in{\color[rgb]{0,0,0}\mathbb{Z}}}\int_{0}^{t}\mathrm{d}\tau\left[r_{13}p_{1,n}(\tau)+r_{31}p_{3,n}(\tau)\right],$ (113)
@@
The quality factors of the TUR $\mathcal{Q}^{\mathrm{TUR}}$ and TB $\mathcal{Q}^{\mathrm{TB}}$ are shown in Fig. 9c (solid and dashed line, respectively) for ${\color[rgb]{0,0,0}e}^{\mathrm{out}}_{B}$ in the antiporter regime. Because $t\mathcal{D}_{t}^{\mathrm{ext}}>\mathrm{var}_{\mathrm{s}}(J_{t}^{\mathrm{SAT}})$ for all $t>0$, the TUR infers a higher fraction of the entropy production, i.e., $\mathcal{Q}^{\mathrm{TUR}}\geq\mathcal{Q}^{\mathrm{TB}}$. Additionally, one may show that taking ${\color[rgb]{0,0,0}e}^{\mathrm{out}}_{B}\to\infty$ leads to the quality factors converging to a finite value [45, 157]. Hence, there is no improvement in the inference of dissipation at higher concentrations of $B$ outside the cell beyond a certain point.
@@ IV.4 Thermodynamic Correlation Bound
where $\tilde{Z}_{xy}(\tau)=Z_{xy}(\tau)\left[z_{x}(\tau)+z_{y}(\tau)-2F(\tau)\right]$ for some time-dependent function $F(\tau)$, $Z_{xy}$ is given by Eq. (76), and $\bm{H}$ is the same matrix also entering Eq. (97). The second key ingredient is the shifted and rescaled density observable
${C}_{t}$ $=\frac{1}{\sqrt{t}}\int_{\tau=0}^{\tau=t}{\color[rgb]{0,0,0}\left[V_{\tau}-\langle{V}_{\tau}\rangle\right]\mathrm{d}{\tau}},$ (116)
where $\langle V_{\tau}\rangle=\sum_{i}V_{i}p_{i}(\tau)$. While the TUR and TB allow the state functions $z_{\tau}$ and ${V}_{\tau}$ to be explicitly time-dependent, the correlation bound requires these to be strictly time-independent. Furthermore, we write, e.g., $\langle z_{\tau}\rangle=\sum_{i}z_{i}p_{i}(\tau)$ to explicitly highlight that the average depends on time through $p_{i}(\tau)$. Additionally, we need to introduce the re-weighted two-point probability
@@
With the shorthand notations $V_{x}^{\Delta}(\tau)=V_{x}-\langle V_{\tau}\rangle$ and $z^{F}_{\tau}=z_{\tau}-\langle F(\tau)\rangle$, the cross term is computed to be
$t\langle B_{t}C_{t}\rangle$ $=\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{x,y,i}{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}V^{\Delta}_{i}(\tau^{\prime})\left(H_{yx}-Z_{xy}(\tau)\left[z_{x}^{F}(\tau)+z_{y}^{F}(\tau)\right]r_{xy}p_{x}(\tau)\right)\left[P(i,\tau^{\prime}|y,\tau)-P(i,\tau^{\prime}|x,\tau)\right],$ (121)
@@
$\sum_{x,y}Z_{xy}(z_{x}^{F}+z_{y}^{F})r_{xy}p_{x}\left[P(i|y)-P(i|x)\right]=\sum_{x,y}z_{x}^{F}(r_{xy}p_{x}-r_{yx}p_{y})\left[P(i|y)-P(i|x)\right].$ (123)
Using time-homogeneity of the semi-group $\bm{P}(\tau^{\prime}|\tau)=\bm{P}(\tau^{\prime}-\tau)\equiv{\rm e}^{\bm{L}(\tau^{\prime}-\tau)}$ which obeys the master equation Eq. (40), i.e. $\partial_{\tau}\bm{P}(\tau)=\bm{L}\bm{P}(\tau)=\bm{P}(\tau)\bm{L}$, and $\partial_{\tau}\bm{P}(\tau^{\prime}-\tau)=-\partial_{\tau^{\prime}}\bm{P}(\tau^{\prime}-\tau)$, the integrand in Eq. (121) simplifies to $2\sum_{x,i}{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}V^{\Delta}_{i}(\tau^{\prime})z^{F}_{x}(\tau)p_{x}(\tau)\partial_{\tau^{\prime}}P(i,\tau^{\prime}|x,\tau)$. Since
@@
$t\langle B_{t}C_{t}\rangle$ $=2\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}\partial_{\tau^{\prime}}\mathrm{cov}({V}_{\tau^{\prime}},z_{\tau}).$ (125)
An integration by parts in $\tau^{\prime}$ with $\partial_{\tau^{\prime}}{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}=\delta(\tau^{\prime}-\tau)$ further yields
@@
$\frac{4\left[\mathrm{cov}({V}_{t},\overline{{z}}_{t})-\mathrm{cov}_{\mathrm{s}}({V},{z})\right]^{2}}{t\mathrm{var}(\overline{V}_{t})}-\mathcal{D}_{t}\leq\frac{\langle A_{t}^{2}\rangle_{\mathrm{s}}}{t}\mathrm{pvar}_{\mathrm{ps}}^{F}({z})\leq\frac{\Delta S_{\mathrm{tot}}(t)}{2t}\mathrm{pvar}_{\mathrm{ps}}^{F}({z}).$ (131)
@@
Example II: 4-State Toy model.—Consider a four-state model $\{1,2,3,4\}$ on a ring, see Fig. 6e. The rates clockwise are unity and the counterclockwise rates are controlled by the parameter $\epsilon$. The steady-state distribution $p_{i}^{\mathrm{s}}=p=1/4$ is independent of $\epsilon$. Suppose that we can only distinguish between a pair of mesostates A and B, i.e., it can only be distinguished whether the system is in either states 1 and 2 (mesostate A) or states 3 and 4 (mesostate B). These we refer to mesostates A (containing $\{1,2\}$) and B (containing $\{3,4\}$), see Fig. 6e.